Find (a) f+g, (b) f-g, (c) f*g, (d) f/g,
f(x)=x^2-4, g(x)=2+x

Answer by palanisamy(496)

GIVEN, f(x)=x^2-4, g(x)=2+x
a) (f+g)x = f(x)+g(x) = x^2-4+2+x = x^2+x-2
b) (f-g)x = f(x)-g(x) = x^2-4-(2+x) = x^2-4-x-2 = x^2-x-6
c) (f*g)x = f(x)*g(x) = (x^2-4)*(2+x) = 2x^2+x^3-8-4x = x^3+2x^2-4x-8
d) (f/g)x = f(x)/g(x) = (x^2-4)/(2+x) = (x+2)(x-2)/(2+x) = x-2